### v1

# Relative consistency of a finite nonclassical theory incorporating ZF and category theory with ZF

Abstract
Recently, in Axioms 10(2): 119 (2021), a nonclassical first-order theory $\mathfrak{T}$ of sets and functions has been introduced as the collection of axioms we have to accept if we want a foundational theory for (all of) mathematics that is not weaker than ZF, that is finitely axiomatized, and that does not have a countable model (if it has a model at all, that is). Here we prove that $\mathfrak{T}$ is relatively consistent with ZF. We conclude that this is an important step towards showing that $\mathfrak{T}$ is an advancement in the foundations of mathematics.
1 Introduction
The most widely accepted foundational theory for mathematics is without a doubt Zermelo-Fraenkel set theory. Its constructive axioms are those of Zermelo’s set theory, introduced in [1], plus the axiom schema of replacement suggested by Fraenkel in [2]. In accordance with common convention, we will use ZFC to denote the full theory, ZF for the full theory minus the axiom of choice (AC), and ZF(C) in statements that are to hold for both ZF and ZFC. It is then a fact that ZF(C) has infinitely many axioms. And as a corollary of the (downward) Löwenheim-Skolem theoren, developed by Löwenheim and Skolem [3, 4], it is a fact that if ZF(C) has a model, then it has a countable model. These two facts can be viewed as unwanted features of ZF(C)—historically, Zermelo himself and also Von Neumann considered the countable model a pathological feature of ZF(C) [5, 6].
Recently a finitely axiomatized nonstandard first-order theory of functions and sets, denoted by $\mathfrak{T}$, has been published as the collection of axioms one has to accept to get rid of these unwanted features [7]. The axioms of $\mathfrak{T}$ consist of some twenty standard first-order axioms, plus one nonstandard first-order axiom. That nonstandard axiom is so powerful that it admits the deduction of the infinite axiom schemata of separation and replacement (translated in the language of $\mathfrak{T}$) from the finitely many axioms of $\mathfrak{T}$. In addition, the (downward) Löwenheim-Skolem theorem does not hold for $\mathfrak{T}$: if $\mathfrak{T}$ has a model, then the model is uncountable. These two results, proven in [7], allow us to say that $\mathfrak{T}$ is stronger than ZF.
The purpose of this paper is to show that $\mathfrak{T}$ is relatively consistent with ZF. For the sake of self-containment of this paper, the nonstandard theory $\mathfrak{T}$ is summarily outlined in the reminder of this introduction. It has to be kept in mind, however, that this summary outline does not replace the formal introduction of $\mathfrak{T}$ in [7].
For starters, the universe of discourse is a category of sets and functions: its objects are sets, its arrows are functions. Importantly, while a function can be an element of a set, it’s not a set: it’s a concatenation of two sets. To explain, let $\{\alpha_{1},\ldots\},\{\alpha_{2},\ldots\},\ldots$ be any sets. Then a concatenation of $n$ sets is a thing
$\{\alpha_{1},\ldots\}\{\alpha_{2},\ldots\}\cdots\{\alpha_{n},\ldots\}$
which for $n>1$ is different from any set $Z$:
$\{\alpha_{1},\ldots\}\{\alpha_{2},\ldots\}\cdots\{\alpha_{n},\ldots\}\neq Z$ (1)
So, very simple, a concatenation of two sets is a thing $\{\alpha,\ldots\}\{\beta,\ldots\}$ which is not a set. (By the same token, an array of two bricks is not one brick.) A function is then a concatenation of two sets—the first is the graph of the function, the second is its domain—with the second one written as a subscript as in $F_{X}$. An ur-function in this framework is a function with a singleton domain and a singleton range. E.g. for Zermelo ordinals 0 and 1, the ur-function from $\{0\}$ to $\{1\}$ that maps 0 to 1 is the thing $\{(0,1)\}_{\{0\}}$ given by
$\{(0,1)\}_{\{0\}}:\{0\}\twoheadrightarrow\{1\}$ (2)
$\{(0,1)\}_{\{0\}}:0\mapsto 1$ (3)
[…]
Figure 1: Graphical representation of the ur-function $\{(0,1)\}_{\{0\}}$. The left oval with the dot inside is a Venn diagram of the set $\{0\}$. The right oval with the dot inside is a Venn diagram of the set $\{1\}$. The arrow then represents the ur-function $\{(0,1)\}_{\{0\}}$ from $\{0\}$ to $\{1\}$ that maps 0 to 1.
That being said, all but one of the axioms of $\mathfrak{T}$ are expressed in a standard first-order language. So, we proceed by listing the terms of the language and its atomic expressions, starting with the terms:
  • constants $\emptyset$ and $\omega$ stand for the empty set and the (infinite) set of all finite Zermelo ordinals;
  • (labeled) Roman variables $x,x_{1},X,y,\ldots$ vary over sets;
  • for a constant set $\mathbf{\hat{X}}$, variables $f_{\mathbf{\hat{X}}},G_{\mathbf{\hat{X}}},\ldots$ vary over functions on the set $\mathbf{\hat{X}}$;
  • the constant $1_{\emptyset}$ stands for the inactive function;
  • for a variable $X$ ranging over sets, composites variables $F_{X},f_{X},h_{X},\ldots$ range over functions on a set $X$;
  • (labeled) Greek variables $\alpha,\alpha_{1},\beta,\ldots$ range over all things (set and functions).
The atomic expressions are the following:
  • $t_{1}=t_{2}$, meaning $t_{1}$ is identical to $t_{2}$;
  • $t_{1}\in t_{2}$, meaning $t_{1}$ is an element of $t_{2}$;
  • $t_{1}:t_{2}\twoheadrightarrow t_{3}$, meaning $t_{1}$ is a surjection from $t_{2}$ onto $t_{3}$;
  • $t_{1}:t_{2}\mapsto t_{3}$, meaning $t_{1}$ maps $t_{2}$ to $t_{3}$.
The standard axioms of $\mathfrak{T}$ can then be divided into set-theoretical axioms and function-theoretical axioms. Without further explanation, the set-theoretical axioms are the following:
$\forall X\forall Y(X=Y\Leftrightarrow\forall\alpha(\alpha\in X\Leftrightarrow\alpha\in Y))$ (4)
$\forall X\forall f_{X}\forall Y(f_{X}\neq Y)$ (5)
$\forall X\forall\alpha\forall\beta(X:\alpha\not\twoheadrightarrow\beta\wedge X:\alpha\not\mapsto\beta)$ (6)
$\exists X(X=\emptyset\wedge\forall\alpha(\alpha\not\in X))$ (7)
$\forall\alpha\forall\beta\exists X\forall\gamma(\gamma\in X\Leftrightarrow\gamma=\alpha\vee\gamma=\beta)$ (8)
$\forall X\exists Y\forall\alpha(\alpha\in Y\Leftrightarrow\exists Z(Z\in X\wedge\alpha\in Z))$ (9)
$\forall X\exists Y\forall\alpha(\alpha\in Y\Leftrightarrow\exists Z(Z\subset X\wedge\alpha=Z))$ (10)
$\forall X\neq\emptyset\exists\alpha(\alpha\in X\wedge\forall\beta(\beta\in\alpha\Rightarrow\beta\not\in X))$ (11)
$\exists X(X=\omega\wedge\emptyset\in\bm{\omega}\wedge\forall\alpha(\alpha\in\bm{\omega}\Rightarrow\alpha^{+}\in\bm{\omega})\wedge\forall\beta\in\bm{\omega}(\not\exists\gamma\in\bm{\omega}(\beta=\gamma^{+})\Leftrightarrow\beta=\emptyset))$ (12)
This last axiom, the infinite set axiom, makes use of the function symbol $(.)^{+}$. This is defined as follows:
$\forall\alpha\forall\beta(\beta=\alpha^{+}\Leftrightarrow\exists X(\beta=X\wedge\forall\gamma(\gamma\in X\Leftrightarrow\gamma=\alpha)))$ (13)
Proceeding, the standard function-theoretic axioms of $\mathfrak{T}$ are the following:
$\forall X\forall f_{X}\forall\alpha(\alpha\not\in f_{X})$ (14)
$\forall X\forall f_{X}(X\neq\emptyset\Rightarrow\exists Y\exists Z(f_{X}:Y\twoheadrightarrow Z\wedge\forall\alpha\in Y\exists!\beta(f_{X}:\alpha\mapsto\beta)))$ (15)
$\forall X\forall f_{X}\forall\alpha(\alpha\neq X\Rightarrow\forall\xi(f_{X}:\alpha\not\twoheadrightarrow\xi))$ (16)
$\forall X\forall f_{X}\forall\alpha\not\in X\forall\beta(f_{X}:\alpha\not\mapsto\beta)$ (17)
$\forall X\forall f_{X}(X\neq\emptyset\Rightarrow\forall\beta(f_{X}:X\twoheadrightarrow\beta\Rightarrow\exists Z(\beta=Z\wedge\forall\gamma(\gamma\in Z\Leftrightarrow\exists\eta\in X(f_{X}:\eta\mapsto\gamma)))))$ (18)
$\forall X\neq\emptyset\forall f_{X}\forall Y(f_{X}:X\twoheadrightarrow Y\Rightarrow\forall\beta\exists Z\forall\alpha(\alpha\in Z\Leftrightarrow\alpha\in X\wedge f_{X}:\alpha\mapsto\beta))$ (19)
$\exists f_{\emptyset}(f_{\emptyset}=1_{\emptyset}\wedge f_{\emptyset}:\emptyset\twoheadrightarrow\emptyset\wedge\forall\alpha\forall\beta(f_{\emptyset}:\alpha\not\mapsto\beta))$ (20)
$\forall\alpha\forall\beta\exists f_{\alpha^{+}}(f_{\alpha^{+}}:\alpha^{+}\twoheadrightarrow\beta^{+}\wedge f_{\alpha^{+}}:\alpha\mapsto\beta)$ (21)
$\forall X\forall f_{X}\forall Y(f_{X}:X\twoheadrightarrow Y\Rightarrow\forall\alpha(f_{X}:f_{X}\not\mapsto\alpha\wedge f_{X}:\alpha\not\mapsto f_{X}))$ (22)
With that, we have listed all standard axioms of $\mathfrak{T}$.
But the power of the theory $\mathfrak{T}$ comes from its Sum Function Axiom, which is formalized in nonstandard logic. We will demonstrate the axiom with a simple example using Venn diagrams, but for a rigorous introduction we refer to the literature [7]. That being said, let $\alpha$, $\beta$, and $\gamma$ be any three different things (set or function), and let $f_{\{\alpha\}}$, $f_{\{\beta\}}$, and $f_{\{\gamma\}}$ be any three ur-functions on the singletons of $\alpha$, $\beta$, and $\gamma$, respectively; let the images of the ur-functions be denoted by $f_{\{\alpha\}}(\alpha)$, $f_{\{\beta\}}(\beta)$ and $f_{\{\gamma\}}(\gamma)$. Then the Sum Function Axiom says that there is a sum function $F_{\{\alpha,\beta,\gamma\}}$ from the set $\{\alpha,\beta,\gamma\}$ onto a set $Y$ that maps each of its elements to its image under the ur-function on the corresponding singleton. We can write this down in a first-order formula with composite variables $g_{Z}$ ranging over functions on the set $Z$:
$\begin{array}[]{l}\forall f_{\{\alpha\}}\forall f_{\{\beta\}}\forall f_{\{\gamma\}}\exists F_{\{\alpha,\beta,\gamma\}}\exists Y:\\ F_{\{\alpha,\beta,\gamma\}}:\{\alpha,\beta,\gamma\}\twoheadrightarrow Y\wedge F_{\{\alpha,\beta,\gamma\}}:\alpha\mapsto f_{\{\alpha\}}(\alpha)\wedge F_{\{\alpha,\beta,\gamma\}}:\beta\mapsto f_{\{\beta\}}(\beta)\wedge F_{\{\alpha,\beta,\gamma\}}:\gamma\mapsto f_{\{\gamma\}}(\gamma)\end{array}$ (23)
This reads as: for any ur-function $f_{\{\alpha\}}$ on the singleton of $\alpha$ and for any ur-function $f_{\{\beta\}}$ on the singleton of $\beta$ and for any ur-function $f_{\{\gamma\}}$ on the singleton of $\gamma$ there is a set $Y$ and a function $F_{\{\alpha,\beta,\gamma\}}$ such that $F_{\{\alpha,\beta,\gamma\}}$ is a surjection from the set $\{\alpha,\beta,\gamma\}$ onto the set $Y$ and such that $F_{\{\alpha,\beta,\gamma\}}$ maps $\alpha$ to $f_{\{\alpha\}}(\alpha)$ and $\beta$ to $f_{\{\beta\}}(\beta)$ and $\gamma$ to $f_{\{\gamma\}}(\gamma)$. Using Venn diagrams, we can give this a graphical representation, see the Fig. 2 below.
Figure 2: Graphical representation of the sum function $F_{\{\alpha,\beta,\gamma\}}$. On the left the three ur-functions $f_{\{\alpha\}}$, $f_{\{\beta\}}$, and $f_{\{\gamma\}}$ are graphically represented as in Fig. 1. On the right, the left oval is a Venn diagram of the set $\{\alpha,\beta,\gamma\}$, the right oval is a Venn diagram of the set $Y$, and the three arrows then represent the postulated sum function $F_{\{\alpha,\beta,\gamma\}}$.
Now we abbreviate formula (23) using a multiple quantifier $(\forall f_{\{x\}})_{x\in\{\alpha,\beta,\gamma\}}$ and a conjunctive operator $\bigwedge\limits_{x\in\{\alpha,\beta,\gamma\}}$ to an equivalent well-formed formula in the language of $\mathfrak{T}$:
$(\forall f_{\{x\}})_{x\in\{\alpha,\beta,\gamma\}}\exists F_{\{\alpha,\beta,\gamma\}}\exists Y\left[F_{\{\alpha,\beta,\gamma\}}:\{\alpha,\beta,\gamma\}\twoheadrightarrow Y\wedge\bigwedge\limits_{x\in\{\alpha,\beta,\gamma\}}F_{\{\alpha,\beta,\gamma\}}:x\mapsto f_{\{x\}}(x)\right]$ (24)
This reads as: for any family of ur-functions $f_{\{x\}}$ indexed in $\{\alpha,\beta,\gamma\}$, there is a set $Y$ and a function $F_{\{\alpha,\beta,\gamma\}}$ such that $F_{\{\alpha,\beta,\gamma\}}$ is a surjection from $\{\alpha,\beta,\gamma\}$ onto the set $Y$ and, conjunctively over all $x\in\{\alpha,\beta,\gamma\}$, $F_{\{\alpha,\beta,\gamma\}}$ maps $x$ to its image under the ur-function $f_{\{x\}}$. It has to be taken that in this simple case the multiple quantifier is equivalent to three standard quantifiers, and the subformula $\bigwedge\limits_{x\in\{\alpha,\beta,\gamma\}}F_{\{\alpha,\beta,\gamma\}}:x\mapsto f_{\{x\}}(x)$ is equivalent to a standard conjunction of three propositions—this may be obvious from comparing formulas (23) and (24). The full Sum Function Axiom, abbreviated SUM-F, now says that formula (24) holds for any (possibly infinite) non-empty set $X$—not just for the set $\{\alpha,\beta,\gamma\}$. So, SUM-F is the axiom
$\forall X(\forall f_{\{x\}})_{x\in X}\exists F_{X}\exists Y\left[F_{X}:X\twoheadrightarrow Y\wedge\bigwedge\limits_{x\in X}F_{X}:x\mapsto f_{\{x\}}(x)\right]$ (25)
This is a new nonstandard mathematical principle; in the general case, the multiple quantifier $(\forall f_{\{x\}})_{x\in X}$ can be equivalent to infinitely many standard quantifiers, and the subformula $\bigwedge\limits_{x\in X}F_{X}:x\mapsto f_{\{x\}}(x)$ can be equivalent to a standard conjunction of infinitely many propositions. The axiomatic system of $\mathfrak{T}$ also contains nonstandard rules of inference, some of which are treated in Sect. 5. For a formal introduction, see [7].
2 Method
We want to prove that $\mathfrak{T}$ is relative consistent with ZF. That is, we want to prove that
$\left[\nvdash_{ZF}\bot\right]\Rightarrow\left[\nvdash_{\mathfrak{T}}\bot\right]$ (26)
[…]
[⊬$\left[\not\vdash_{ZF}\bot\right]\Rightarrow\left[\not\Vvdash_{\mathfrak{T}}\bot\right]$ (27)
We will prove this proposition in three steps:
  1. (i)
    we first define the standard first-order theory $\mathfrak{T}^{\infty}_{0}$ as the theory $\mathfrak{T}$ minus the nonstandard Sum Function Axiom but extended with a collection of standard first-order theorems of $\mathfrak{T}$ (Def. 3.1), and we prove that $\mathfrak{T}^{\infty}_{0}$ is relative consistent with ZF as in
    $\left[\nvdash_{ZF}\bot\right]\Rightarrow\left[\nvdash_{\mathfrak{T}^{\infty}_{0}}\bot\right]$ (28)
  2. (ii)
    we then show that $\mathfrak{T}^{\infty}_{0}$ in the nonstandard logical framework of $\mathfrak{T}$ is relatively consistent with $\mathfrak{T}^{\infty}_{0}$ in the framework of standard first-order logic, as in
    [⊬$\left[\not\vdash_{\mathfrak{T}^{\infty}_{0}}\bot\right]\Rightarrow\left[\not\Vvdash_{\mathfrak{T}^{\infty}_{0}}\bot\right]$ (29)
  3. (iii)
    finally, we prove that $\mathfrak{T}$ is relatively consistent with $\mathfrak{T}^{\infty}_{0}$ as in
    [⊪̸$\not\Vvdash_{\mathfrak{T}^{\infty}_{0}}\bot\right]\Rightarrow\left[\not\Vvdash_{\mathfrak{T}}\bot\right][ ⊪$ (30)
By transitivity of relative consistency, we may then infer that $\mathfrak{T}$ is relatively consistent with ZF as in Eq. (27). The next three sections treat the above three steps one by one.
3 Relative consistency of $\mathfrak{T}^{\infty}_{0}$ with ZF
For this first step, we start from Set Matrix Theory (SMT), introduced in [8] as a generalisation of ZF in which finite matrices with set-valued entries are treated as objects sui generis; an outline of SMT is given in the Appendix. In [9] it has been proven that SMT is relative consistent with ZF. We then proceed as follows:
  1. (i)
    we define STM1×2 as a restriction of SMT in which only set matrices of the type $[f\ \ X]$ occur where $f$ is the graph of a function on a set $X$ (Def. 3.2);
  2. (ii)
    we show that that SMT1×2 is relatively consistent with SMT (Prop. 3.3);
  3. (iii)
[…]
  4. (iv)
[…]
  5. (v)
[…]
We may then infer that $\mathfrak{T}^{\infty}_{0}$ is relatively consistent with SMT. By transitivity of relative consistency, we may then infer that $\mathfrak{T}^{\infty}_{0}$ is relatively consistent with ZF.
For starters, recall from [7] that the theorem schemata of separation and replacement of $\mathfrak{T}$ are the schemata
$\forall X\exists Y\forall\alpha:\alpha\in Y\Leftrightarrow\alpha\in X\wedge\Phi(\alpha)$ (31)
$\forall X:\forall\alpha\in X\exists!\beta\Psi(\alpha,\beta)\Rightarrow\exists Z\forall\gamma(\gamma\in Z\Leftrightarrow\exists\xi:\xi\in X\wedge\Psi(\xi,\gamma))$ (32)
with subformulas $\Phi$ and $\Psi$ interpreted as usual. In addition, the following is a theorem of $\mathfrak{T}$:
$\forall f\forall X\forall Y:f\in Y^{X}\Rightarrow\exists h_{X}\forall\alpha\forall\beta(h_{X}:\alpha\mapsto\beta\Leftrightarrow(\alpha,\beta)\in f)$ (33)
where $Y^{X}$ is the set of all graphs of functions from $X$ to $Y$; the proof is omitted.
Definition 3.1.
Let $\Gamma$ be the collection of formulas in the language of $\mathfrak{T}$ made up of all theorems in the schemata (31) and (32) plus theorem (33). Let $\mathfrak{T}[\Gamma]$ be the supertheory of $\mathfrak{T}$ obtained by adding the formulas of $\Gamma$ as axioms to $\mathfrak{T}$. Then $\mathfrak{T}^{\infty}_{0}$ is the subtheory of $\mathfrak{T}[\Gamma]$ obtained by removing the nonstandard sum function axiom from the axioms of $\mathfrak{T}[\Gamma]$, and by removing the nonstandard deduction rules. $\blacksquare$
The theory $\mathfrak{T}^{\infty}_{0}$ has, thus, infinitely many axioms (as indicated by the superscript $\infty$) but zero nonstandard axioms (as indicated by the subscript 00). The axioms of $\mathfrak{T}^{\infty}_{0}$ are, thus, the standard set-theoretical and function-theoretical axioms of $\mathfrak{T}$ listed in the introduction, plus all formulas in $\Gamma$ as listed above.
Definition 3.2.
Let the theory SMT1×2 be the subtheory of SMT obtained by
  1. (i)
    removing all function symbols $f_{m\times n}$ from the language, except $f_{1\times 2}$—note that this eliminates the reduction axiom and the omission axiom scheme, given by Eqs. (59) and (60) in the Appendix Appendix: outline of Set Matrix Theory—so that Greek variables $\alpha,\beta,\ldots$ now range over sets and $1\times 2$ matrices;
  2. (ii)
    replacing the set matrix axiom schema, given by Eq. (58) in Appendix Appendix: outline of Set Matrix Theory, by
    $\left(\forall f\forall X\forall Y:f\in Y^{X}\Rightarrow\exists\alpha(\alpha=[f\ \ X])\right)\wedge\left(\forall\alpha:\exists\beta\exists\gamma(\alpha=[\beta\ \ \gamma])\Rightarrow\exists X\exists f\exists Y(\alpha=[f\ \ X]\wedge f\in Y^{X})\right)$ (34)
  3. (iii)
    removing the set of set matrices axiom scheme, given by Eq. (72) in Appendix Appendix: outline of Set Matrix Theory
$\blacksquare$
In the framework of SMT we can construct set matrices $\left[\begin{array}[]{ccc}\alpha_{11}&\ldots&\alpha_{1n}\\ \vdots&&\vdots\\ \alpha_{m1}&\ldots&\alpha_{mn}\end{array}\right]$ of any size $m\times n$ and with any entries $\alpha_{ij}$. In the framework of SMT1×2, however, we only want to allow $1\times 2$ set matrices $[f\ \ X]$ with set-valued entries $f$ and $X$ satisfying the condition that $f$ is the graph of a function on $X$. This restriction forces us to remove the set of set matrices axiom scheme, because that would introduce the existence of $1\times 2$ set matrices $[\alpha\ \ \beta]$ violating this condition. Altogether, we then end up with Def. 3.2.
As a side note, in the framework of SMT, a cartesian product of sets $A$ and $B$ is just the set of $1\times 2$ matrices $[\alpha\ \ \beta]$ with $\alpha\in A$ and $\beta\in B$, which can be constructed from the set of set matrices $M_{1\times 2}(A\cup B)$ and a separation axiom. But in the framework of SMT1×2 we can no longer construct the cartesian product $A\times B$ that way, because we removed the set of set matrices axiom scheme. So in order to be able construct the cartesian product $A\times B$ in the framework of SMT1×2, we are forced to define a two-tuple $\langle\alpha,\beta\rangle$ as a set—we can use any one of the usual definitions, e.g. $\langle\alpha,\beta\rangle=\{\alpha,\{\alpha,\beta\}\}$.
Proposition 3.3.
SMT1×2 is relatively consistent with SMT. $\blacksquare$
Proof:
The proof is omitted. $\blacksquare$
Definition 3.4.
[…]
  1. (i)
    a definition of a ternary relation $t_{1}:t_{2}\twoheadrightarrow t_{3}$:
    $\forall\alpha\forall\beta\forall\gamma\left(\alpha:\beta\twoheadrightarrow\gamma\Leftrightarrow\exists f\exists X\exists Y:\alpha=[f\ \ X]\wedge\beta=X\wedge\gamma=Y\wedge\forall\mu\in Y\exists\nu\in X(\langle\nu,\mu\rangle\in f)\right)$ (35)
  2. (ii)
    a definition of a ternary relation $t_{1}:t_{2}\mapsto t_{3}$:
    $\forall\alpha\forall\beta\forall\gamma\left(\alpha:\beta\mapsto\gamma\Leftrightarrow\exists X\exists f:\alpha=[f\ \ X]\wedge\langle\beta,\gamma\rangle\in f\right)$ (36)
  3. (iii)
    an axiom for the constant $\omega$ representing the set of all Zermelo ordinals:
    $\forall x:x=\omega\Leftrightarrow\emptyset\in x\wedge\forall\alpha(\alpha\in x\Rightarrow\alpha^{+}\in x)\wedge\forall\beta\in x(\not\exists\gamma\in x(\beta=\gamma^{+})\Leftrightarrow\beta=\emptyset))$ (37)
  4. (iv)
[…]
    $\forall\alpha:\alpha=[\emptyset\ \ \emptyset]\Leftrightarrow\alpha:\emptyset\twoheadrightarrow\emptyset\wedge\forall\beta\forall\gamma(\alpha:\beta\not\mapsto\gamma)$ (38)
$\blacksquare$
[…]
  1. (i)
    $\alpha$ is the $1\times 2$ set matrix $[f\ \ X]$—note that by Eq. (34), this implies that $f$ is the graph of a function on $X$;
  2. (ii)
    $\beta$ is the set $X$;
  3. (iii)
    $\gamma$ is the set $Y$ and for every element $\mu$ in $Y$ there is an element $\nu$ in $X$ such that $\langle\nu,\mu\rangle$ is in $f$.
Furthermore, the object $\alpha$ maps $\beta$ to $\gamma$ if and only if there is a set $X$ and a graph $f$ of a function on $X$ such that $\alpha$ is the $1\times 2$ set matrix $[f\ \ X]$ and the two-tuple $\langle\beta,\gamma\rangle$ is in $f$.
Proposition 3.5.
[…]
Proof:
[…]
Firstly, Eqs. (35) and (36) are defining axioms of the ternary predicates that generate the atomic expressions $t_{1}:t_{2}\twoheadrightarrow t_{3}$ and $t_{1}:t_{2}\mapsto t_{3}$. The extension of SMT1×2 by Eqs. (35) and (36) is therefore an extension by definitions: this is a conservative extension [10].
[…]
Proposition 3.6.
[…]
Proof:
[…]
  • $\tau:\emptyset\mapsto\emptyset$ , $\tau:\omega\mapsto\omega$ , $\tau:1_{\emptyset}\mapsto[\emptyset\ \ \emptyset]$;
  • a simple variable $x$ ranging over sets is mapped by $\tau$ to a simple variable $x$ ranging over sets;
  • a simple variable $\alpha$ ranging over all things (sets and functions) is mapped by $\tau$ to a simple variable $\alpha$ ranging over all things (sets and set matrices);
  • for any constant $\bm{\hat{X}}$ of $L_{\mathfrak{T}^{\infty}_{0}}$ referring to an individual set, a composite term $f_{\bm{\hat{X}}}$ ranging over functions on the set $\bm{\hat{X}}$ is mapped by $\tau$ to a $1\times 2$ matrix $[f\ \ \bm{\hat{X}}]$ which effectively ranges over functions on the set $\bm{\hat{X}}$;
  • a composite variable $f_{X}$ ranging over functions on a set $X$ is mapped by $\tau$ to a $1\times 2$ matrix $[f\ \ X]$ which effectively ranges over functions on a set $X$;
  • atomic formulas $t_{1}\in t_{2}$, $t_{1}=t_{2}$, $t_{1}:t_{2}\twoheadrightarrow t_{3}$, and $t_{1}:t_{2}\mapsto t_{3}$ are mapped by $\tau$ to respectively formulas $\tau(t_{1})\in\tau(t_{2})$, $\tau(t_{1})=\tau(t_{2})$, $\tau(t_{1}):\tau(t_{2})\twoheadrightarrow\tau(t_{3})$, and $\tau(t_{1}):\tau(t_{2})\mapsto\tau(t_{3})$;
  • formulas $\neg\Phi$ and $\Phi\ldots\Psi$ with logical connective $\ldots$’ are mapped by $\tau$ to formulas $\neg\tau(\Phi)$ and $\tau(\Phi)\ldots\tau(\Psi)$;
  • for any simple variable $\alpha$ ranging over all things, a formula $\forall\alpha:\Phi$ is mapped by $\tau$ to a formula $\forall\alpha:\tau(\Phi)$;
  • for any simple variable $X$ ranging over sets, a formula $\forall X:\Phi$ is mapped by $\tau$ to a formula $\forall X:\tau(\Phi)$;
  • for any constant $\bm{\hat{X}}$ of $L_{\mathfrak{T}^{\infty}_{0}}$ referring to an individual set, a formula $\forall f_{\bm{\hat{X}}}:\Phi$ or $\exists f_{\bm{\hat{X}}}:\Phi$ is mapped by $\tau$ to a formula $\forall f:\exists\alpha\left(\alpha=[f\ \ \bm{\hat{X}}]\right)\Rightarrow\tau(\Phi)$ or $\exists f:\exists\alpha\left(\alpha=[f\ \ \bm{\hat{X}}]\right)\wedge\tau(\Phi)$;
  • for any simple variable $X$ ranging over sets, a formula $\forall f_{X}:\Phi$ or $\exists f_{X}:\Phi$ in the scope of the quantifier $\forall X$ is mapped by $\tau$ to a formula $\forall f:\exists\alpha\left(\alpha=[f\ \ X]\right)\Rightarrow\tau(\Phi)$ or $\exists f:\exists\alpha\left(\alpha=[f\ \ X]\right)\wedge\tau(\Phi)$
[…]
$\forall f\forall X\forall Y:f\in Y^{X}\Rightarrow\exists h(\exists\gamma(\gamma=[h\ \ X])\wedge\forall\alpha\forall\beta([h\ \ X]:\alpha\mapsto\beta\Leftrightarrow(\alpha,\beta)\in f)$ (39)
[…]
$f\in Y^{X}$ (40)
for arbitrary $f,X,Y$. From assumption (40) and axiom (34) we infer
$\exists\gamma:\gamma=[f\ \ X]$ (41)
From Eq. (41) and axiom (36) we then infer
$\forall\alpha\forall\beta([f\ \ X]:\alpha\mapsto\beta\Leftrightarrow(\alpha,\beta)\in f)$ (42)
From here we deduce
$f\in Y^{X}\Rightarrow\exists\gamma:\gamma=[f\ \ X])\wedge\forall\alpha\forall\beta([f\ \ X]:\alpha\mapsto\beta\Leftrightarrow(\alpha,\beta)\in f)$ (43)
We then obtain Eq. (39) by introducing existential and universal quantifiers. $\blacksquare$
Corollary 3.7.
$\mathfrak{T}^{\infty}_{0}$ is relatively consistent with ZF. $\blacksquare$
4 Extending the framework with nonclassical rules of inference
To prove Eq. (29), we add the nonstandard rules of inference to the standard logical framework of $\mathfrak{T}^{\infty}_{0}$, and show that this doesn’t introduce any inconsistencies. For that matter, it suffices to prove the following proposition.
Proposition 4.1.
Let $\Sigma$ be a collection of standard first-order formulas in the language $L(\mathfrak{T})$ of the nonstandard theory $\mathfrak{T}$. Then, if no contradiction can be derived from $\Sigma$ using the standard rules of inference of first-order logic, no contradiction can be derived from $\Sigma$ using the standard and nonstandard rules of inference of the axiomatic system of $\mathfrak{T}$:
$\left[\Sigma\not\vdash\bot\right]\Rightarrow\left[\Sigma\not\Vvdash\bot\right]$̸ ⊥ ] (44)
$\blacksquare$
Proof:
Let $\Sigma_{1}$ be the collection of formulas made up of the formulas $\phi\in\Sigma$ and the formulas that can be derived from formulas in $\Sigma$ by once applying the rules of inference of standard first-order logic. Let $\Sigma_{2}$ be the collection of formulas made up of the formulas $\phi\in\Sigma_{1}$ and the formulas that can be derived from formulas in $\Sigma_{1}$ by once applying the rules of inference of standard first-order logic, and so forth. Let $\Sigma_{\omega}$ be the collection of formulas such that $\phi\in\Sigma_{\omega}$ if and only if there is an integer $n$ for which $\phi\in\Sigma_{n}$. We now assume
$\Sigma\not\vdash\bot$ (45)
So, there is no formula $\phi$ such that $\phi\in\Sigma_{\omega}\wedge\neg\phi\in\Sigma_{\omega}$. Now there are just a handful of nonstandard rules of inference in the axiomatic system of $\mathfrak{T}$, so we start by treating the first one: introduction of a conjunctive operator $\bigwedge\limits_{\alpha\in\mathbf{\hat{X}}}$ for a constant $\mathbf{\hat{X}}\in L(\mathfrak{T})$, given by
$\left\{\mathbf{\hat{F}}_{\mathbf{\hat{X}}}:\alpha\mapsto\mathbf{\hat{f}}_{\{\alpha\}}(\alpha)\right\}_{\alpha\in\mathbf{\hat{X}}}\Vvdash\bigwedge\limits_{\alpha\in\mathbf{\hat{X}}}\mathbf{\hat{F}}_{\mathbf{\hat{X}}}:\alpha\mapsto\mathbf{\hat{f}}_{\{\alpha\}}(\alpha)$ (46)
Let’s call this ‘rule $I_{1}$’. Now let $\Upsilon_{1}$ be the collection of formulas made up of the formulas $\phi\in\Sigma_{\omega}$ and the formulas that can be derived from formulas in $\Sigma_{\omega}$ by once applying rule $I_{1}$. Since rule $I_{1}$ only applies to collections of formulas of the type $t_{1}:t_{2}\mapsto t_{3}$, this only adds formulas of the type $\bigwedge\limits_{\alpha\in\mathbf{\hat{X}}}\mathbf{\hat{F}}_{\mathbf{\hat{X}}}:\alpha\mapsto\mathbf{\hat{f}}_{\{\alpha\}}(\alpha)$ to $\Sigma_{\omega}$. But such a formula cannot be the negation of a formula $\phi\in\Sigma_{\omega}$, because it contains a new logical symbol (conjunctive operator) that does not occur in any of the formulas that are in $\Sigma_{\omega}$. Ergo, there is no formula $\phi$ such that $\phi\in\Upsilon_{1}\wedge\neg\phi\in\Upsilon_{1}$.
Proceeding, let $\Upsilon_{2}$ be the collection of formulas made up of the formulas $\phi\in\Upsilon_{1}$ and the formulas that can be derived from formulas in $\Upsilon_{1}$ by once applying one of the standard rules or rule $I_{1}$, and so forth; let $\Upsilon_{\omega}$ be the collection of formulas such that $\phi\in\Upsilon_{\omega}$ if and only if there is an integer $n$ for which $\phi\in\Upsilon_{n}$. But rule $I_{1}$ only applies to collections of atomic formulas of the type $t_{1}:t_{2}\mapsto t_{3}$—it doesn’t apply to any collection of propositions as in infinitary logic—and that means that we cannot create new nonstandard formulas of the type $\bigwedge\limits_{\alpha\in\mathbf{\hat{X}}}\mathbf{\hat{F}}_{\mathbf{\hat{X}}}:\alpha\mapsto\mathbf{\hat{f}}_{\{\alpha\}}(\alpha)$ when we go from $\Upsilon_{n}$ to $\Upsilon_{n+1}$. And that means that $\Upsilon_{\omega}$ consists of all the formulas that can be derived from the formulas $\phi\in\Upsilon_{1}$ by standard propositional logic. Ergo, given the consistency of standard propositional logic, there is no formula $\phi$ such that $\phi\in\Upsilon_{\omega}\wedge\neg\phi\in\Upsilon_{\omega}$.
We now add the second nonstandard rule of inference: elimination of a conjunctive operator $\bigwedge\limits_{\alpha\in\mathbf{\hat{X}}}$ for a constants $\mathbf{\alpha},\mathbf{\hat{X}}\in L(\mathfrak{T})$, with $\mathbf{\alpha}\in\mathbf{\hat{X}}$, given by
$\bigwedge\limits_{\alpha\in\mathbf{\hat{X}}}\mathbf{\hat{F}}_{\mathbf{\hat{X}}}:\alpha\mapsto\mathbf{\hat{f}}_{\{\alpha\}}(\alpha)\Vvdash\mathbf{\hat{F}}_{\mathbf{\hat{X}}}:\mathbf{\alpha}\mapsto\mathbf{\hat{f}}_{\{\mathbf{\alpha}\}}(\mathbf{\alpha})$ (47)
Let’s call this ‘rule $I_{2}$’. Let $\Phi_{1}$ then be the collection of formulas made up of the formulas $\phi\in\Upsilon_{\omega}$ and the formulas that can be derived from formulas in $\Upsilon_{\omega}$ by once applying rule $I_{2}$. But by applying rule $I_{2}$, we can only derive formulas $\phi$ of the type $t_{1}:t_{2}\mapsto t_{3}$ that were already in $\Sigma_{\omega}$. So, if we define $\Phi_{n+1}$ as the collection of formulas made up of the formulas $\phi\in\Phi_{n}$ and the formulas that can be derived from formulas in $\Phi_{n}$ by once applying one of the standard rules or one of the nonstandard rules $I_{1}$ or $I_{2}$, and we define $\Phi_{\omega}$ as the collection of formulas such that $\phi\in\Phi_{\omega}$ if and only if there is an integer $n$ for which $\phi\in\Phi_{n}$, then there is no formula $\phi$ such that $\phi\in\Phi_{\omega}\wedge\neg\phi\in\Phi_{\omega}$.
The other nonstandard inference rules (introduction of a multiple quantifier, elimination of a multiple quantifier, introduction of a single quantifier for a nonstandard formula, elimination of a single quantifier for a nonstandard formula) can be treated the same way—this is quite laborious, but straightforward. This gives a collection of formulas $\Omega_{\omega}$ that can be derived from the collection $\Sigma$ by applying the standard and nonstandard rules of inference of the axiomatic system of $\mathfrak{T}$: there is then no formula $\phi$ such that $\phi\in\Omega_{\omega}\wedge\neg\phi\in\Omega_{\omega}$. That means that $\Sigma$ is also consistent in the axiomatic system of $\mathfrak{T}$: we then have $\Sigma\not\Vvdash\bot$̸ ⊥. Prop. 4.1 then follows immediately. $\blacksquare$
Corollary 4.2.
$\left[\mathfrak{T}^{\infty}_{0}\not\vdash\bot\right]\Rightarrow\left[\mathfrak{T}^{\infty}_{0}\not\Vvdash\bot\right]$̸ ⊥ ] (48)
$\blacksquare$
5 Relative consistency of $\mathfrak{T}$ with $\mathfrak{T}^{\infty}_{0}$
Our proof method is as follows. We assume $\mathfrak{T}^{\infty}_{0}\not\Vvdash\bot$̸ ⊥. That is, we assume that $\mathfrak{T}^{\infty}_{0}$ is consistent. If we then prove that the negation of SUM-F cannot be derived from $\mathfrak{T}^{\infty}_{0}$ (Prop. 5.1) we may derive that $\mathfrak{T}^{\infty}_{0}\cup\{{\rm SUM-F}\}$ is consistent. But $\mathfrak{T}^{\infty}_{0}\cup\{{\rm SUM-F}\}$ is just the theory $\mathfrak{T}[\Gamma]$ of Def. 3.1, which of course is equivalent to $\mathfrak{T}$. Ergo, if we prove Prop. 5.1, then we may conclude that $\mathfrak{T}$ is relatively consistent with $\mathfrak{T}^{\infty}_{0}$ as in Eq. (30).
Proposition 5.1.
The negation of SUM-F cannot be derived from $\mathfrak{T}^{\infty}_{0}$:
$\mathfrak{T}^{\infty}_{0}\not\Vvdash\neg{\rm SUM-F}$ (49)
$\blacksquare$
Proof:
SUM-F is the nonstandard axiom of $\mathfrak{T}$ given by Eq. (25); its negation is the formula
$\exists X(\exists f_{\{x\}})_{x\in X}:\neg\exists F_{X}\exists Y\left(F_{X}:X\twoheadrightarrow Y\right)\wedge\bigwedge\limits_{x\in X}F_{X}:x\mapsto f_{\{x\}}(x)$ (50)
So, we want to prove that it cannot be deduced from $\mathfrak{T}^{\infty}_{0}$ that there is a set $X$ and a family of ur-functions $f_{\{x\}}$ indexed in $X$ such that there is no function $F_{X}$ and no set $Y$ such that $F_{X}$ is a surjection from $X$ onto $Y$ and, conjunctively over $x\in X$, $F_{X}$ maps $x$ to its image under the corresponding ur-function $f_{\{x\}}$. We proceed by reductio ad absurdum: we assume Eq. (50). Applying Rule-C, i.e. the rule $\exists x\Phi(x)\vdash\Phi(\mathbf{x})$ where $\mathbf{x}$ is a constant for which $\Psi$ is true, that means that there is some individual set $\mathbf{\hat{X}}$ and a family of ur-functions $(\mathbf{\hat{f}}_{\{x\}})_{x\in\mathbf{\hat{X}}}$ such that
$\neg\exists F_{\mathbf{\hat{X}}}\exists Y\left(F_{\mathbf{\hat{X}}}:\mathbf{\hat{X}}\twoheadrightarrow Y\right)\wedge\bigwedge\limits_{x\in\mathbf{\hat{X}}}F_{\mathbf{\hat{X}}}:x\mapsto\mathbf{\hat{f}}_{\{x\}}(x)$ (51)
However, we have that
$\forall x\in\mathbf{\hat{X}}\exists!\alpha:\alpha=\mathbf{\hat{f}}_{\{x\}}(x)$ (52)
$\forall x\in\mathbf{\hat{X}}\exists!\beta:\beta=(x,\mathbf{\hat{f}}_{\{x\}}(x))$ (53)
By the axiom schema of substitution (32) of $\mathfrak{T}^{\infty}_{0}$, we thus obtain that
$\exists Y:Y=\{\mathbf{\hat{f}}_{\{x\}}(x)\ |\ x\in\mathbf{\hat{X}}\}$ (54)
$\exists G:G=\{(x,\mathbf{\hat{f}}_{\{x\}}(x))\ |\ x\in\mathbf{\hat{X}}\}$ (55)
meaning that there are sets $Y$ and $G$ such that $Y$ is the set of all images of elements of $\mathbf{\hat{X}}$ under the ur-functions $(\mathbf{\hat{f}}_{\{x\}})_{x\in\mathbf{\hat{X}}}$ and $G$ is the set that contains a two-tuple $(x,\mathbf{\hat{f}}_{\{x\}}(x))$ for every $x\in\mathbf{\hat{X}}$. But then $G$ is the graph of a function on $\mathbf{\hat{X}}$, so by axiom (33) of $\mathfrak{T}^{\infty}_{0}$ we have
$\exists F_{\mathbf{\hat{X}}}\forall x\in\mathbf{\hat{X}}\left(F_{\mathbf{\hat{X}}}:x\mapsto\mathbf{\hat{f}}_{\{x\}}(x)\right)$ (56)
Using the nonstandard deduction rule $\left\{\mathbf{\hat{F}}_{\mathbf{\hat{X}}}:x\mapsto\mathbf{\hat{f}}_{\{x\}}(x)\right\}_{x\in\mathbf{\hat{X}}}\Vvdash\bigwedge\limits_{x\in\mathbf{\hat{X}}}\mathbf{\hat{F}}_{\mathbf{\hat{X}}}:x\mapsto\mathbf{\hat{f}}_{\{x\}}(x)$ we derive from here that
$\exists F_{\mathbf{\hat{X}}}\exists Y\left(F_{\mathbf{\hat{X}}}:\mathbf{\hat{X}}\twoheadrightarrow Y\wedge\bigwedge\limits_{x\in\mathbf{\hat{X}}}F_{\mathbf{\hat{X}}}:x\mapsto\mathbf{\hat{f}}_{\{x\}}(x)\right)$ (57)
which contradicts (51). Ergo, our assumption (50) cannot be true. That proves Prop. 5.1. $\blacksquare$
6 Conclusion
Summarizing, we have proven that $\mathfrak{T}$ is relatively consistent with ZF, as in $\left[\nvdash_{ZF}\bot\right]\Rightarrow\left[\not\Vvdash_{\mathfrak{T}}\bot\right]$. Note that this is a relation between two theories in different logical frameworks! Given that $\mathfrak{T}$ has been proven to be stronger than ZF, we therefore tentatively conclude that we have made an important step towards proving that $\mathfrak{T}$ is an advancement in the foundations of mathematics.
Of course, one may of doubt that mainstream mathematicians will ever use nonclassical axioms such as the sum function axiom in everyday mathematical practice. The point, however, is that they don’t have to. Since ZF can be derived from $\mathfrak{T}$, in everyday mathematical practice it suffices to use the standard first-order theorems of $\mathfrak{T}$: the nonclassical part of $\mathfrak{T}$ stays in the background. The bottom line is that if we start “doing mathematics” in the framework of $\mathfrak{T}$, we don’t have to worry about a countable model—it doesn’t exist.
Appendix: outline of Set Matrix Theory
Informally, SMT is a generalisation of ZF: in the universe of SMT not only sets exist, but also set matrices that are constructed from a finite number of sets—these set matrices are urelements in the sense that they are not sets but can be elements of sets. The language $L_{SMT}$ of SMT is the language $L_{ZF}$ of ZF, with the constant $\emptyset$ and with (labeled) Roman variables $x,y,\ldots$ varying over sets, extended with
  1. (i)
    (labeled) Greek variables $\alpha,\beta,\ldots$ varying over set matrices;
  2. (ii)
    for each pair of positive integers $m$ and $n$, an $mn$-ary function symbols $f_{m\times n}$, with matrix notation $\left[\begin{array}[]{ccc}t_{11}&\ldots&t_{1n}\\ \vdots&&\vdots\\ t_{m1}&\ldots&t_{mn}\end{array}\right]$ for $f_{m\times n}(t_{11},\ldots,t_{mn})$.
A constructive set matrix axiom schema states that given any $m\cdot n$ set matrices $\alpha_{11},\ldots,\alpha_{mn}$, there exists an $m\times n$ set matrix $\beta$ whose entries are the $\alpha_{ij}$’s:
$\forall\alpha_{11}\cdots\forall\alpha_{mn}\exists\beta:\beta=\left[\begin{array}[]{ccc}\alpha_{11}&\ldots&\alpha_{1n}\\ \vdots&&\vdots\\ \alpha_{m1}&\ldots&\alpha_{mn}\end{array}\right]$ (58)
The reduction axiom of SMT identifies a $1\times 1$ set matrix with its entry:
$\forall x:[x]=x$ (59)
This way, quantification over all set matrices implies quantification over all sets.
Likewise, the omission axiom schema identifies a $1\times 1$ set matrix whose entry is an $m\times n$ set matrix with its entry:
$\forall\alpha_{11}\cdots\forall\alpha_{mn}:\left[\left[\begin{array}[]{ccc}\alpha_{11}&\ldots&\alpha_{1n}\\ \vdots&&\vdots\\ \alpha_{m1}&\ldots&\alpha_{mn}\end{array}\right]\right]=\left[\begin{array}[]{ccc}\alpha_{11}&\ldots&\alpha_{1n}\\ \vdots&&\vdots\\ \alpha_{m1}&\ldots&\alpha_{mn}\end{array}\right]$ (60)
The division axiom schema of SMT then states that set matrices of different sizes are not identical; this schema includes axioms stating that a set matrix of any size other than $1\times 1$ is not a set:
$\forall x\forall\alpha_{11}\cdots\forall\alpha_{mn}:x\neq\left[\begin{array}[]{ccc}\alpha_{11}&\ldots&\alpha_{1n}\\ \vdots&&\vdots\\ \alpha_{m1}&\ldots&\alpha_{mn}\end{array}\right]\ \ \ m\cdot n\geq 2$ (61)
And the epsilon axiom schema states that an $m\times n$ set matrix with $m\cdot n\geq 2$ has no elements in the sense of the $\in$-relation:
$\forall\alpha_{11}\cdots\forall\alpha_{mn}\forall\beta:\beta\not\in\left[\begin{array}[]{ccc}\alpha_{11}&\ldots&\alpha_{1n}\\ \vdots&&\vdots\\ \alpha_{m1}&\ldots&\alpha_{mn}\end{array}\right]\ \ m\cdot n\geq 2$ (62)
The extensionality axiom schema for set matrices then states that two $m\times n$ set matrices are equal if their entries are equal:
$\forall\alpha_{11}\cdots\forall\alpha_{mn}\forall\beta_{11}\cdots\forall\beta_{mn}:\left[\begin{array}[]{ccc}\alpha_{11}&\ldots&\alpha_{1n}\\ \vdots&&\vdots\\ \alpha_{m1}&\ldots&\alpha_{mn}\end{array}\right]=\left[\begin{array}[]{ccc}\beta_{11}&\ldots&\beta_{1n}\\ \vdots&&\vdots\\ \beta_{m1}&\ldots&\beta_{mn}\end{array}\right]\Leftrightarrow\alpha_{11}=\beta_{11}\wedge\ldots\wedge\alpha_{mn}=\beta_{mn}$ (63)
As to the set-theoretical axioms of SMT, these are generalizations of the axioms of ZF, since set matrices that are not sets can still be elements of sets. E.g. the generalized extensionality axiom for sets becomes
$\forall x\forall y:x=y\Leftrightarrow\forall\alpha(\alpha\in x\Leftrightarrow\alpha\in y)$ (64)
Without further explanation, the generalizations of the constructive axioms of ZF are the following:
$\exists x:x=\emptyset\wedge\forall\alpha(\alpha\not\in x)$ (65)
$\forall x\exists y\forall\alpha:\alpha\in y\Leftrightarrow\alpha\in x\wedge\Phi(\alpha)$ (66)
$\forall\alpha\forall\beta\exists x\forall\gamma:\gamma\in x\Leftrightarrow\gamma=\alpha\vee\gamma=\beta$ (67)
$\forall x:\forall\alpha(\alpha\in x\Rightarrow\exists u(u=\alpha))\Rightarrow\exists y\forall\beta(\beta\in y\Leftrightarrow\exists z(z\in x\wedge\beta\in z))$ (68)
$\forall x\exists y\forall\alpha:\alpha\in y\Leftrightarrow\exists u(u=\alpha\wedge u\subseteq x)$ (69)
$\exists x:\emptyset\in x\wedge\forall y(y\in x\Rightarrow\{y\}\in x)$ (70)
$\forall x:\forall\alpha\in x\exists!\beta\Phi(\alpha,\beta)\Rightarrow\exists y\forall\zeta(\zeta\in y\Leftrightarrow\exists\gamma:\gamma\in x\wedge\Phi(\gamma,\zeta))$ (71)
In addition, SMT has the set of set matrices axiom schema:
$\forall x\exists y\forall\alpha:\alpha\in y\Leftrightarrow\exists\beta_{11}\ldots\exists\beta_{mn}(\alpha=\left[\begin{array}[]{ccc}\beta_{11}&\ldots&\beta_{1n}\\ \vdots&&\vdots\\ \beta_{m1}&\ldots&\beta_{mn}\end{array}\right]\wedge\beta_{11}\in x\wedge\ldots\wedge\beta_{mn}\in x)$ (72)
These are all the constructive axioms of SMT.
Author contributions
A.R.D. Mathias suggested the proof method. M.J.T.F. Cabbolet developed the proof and wrote the paper.
Acknowledgement
The authors would like to thank Harrie de Swart (Tilburg University) for his useful comments.